OpenStudy (anonymous):

A box contains four letter tiles with the letters M, A, T, and H. A spinner has five equal sections numbered 1 through 5. What is the probability of drawing a tile labeled T from the box and spinning an odd number on the spinner?

OpenStudy (aaronq):

so in the box, you have 4 options, and 1 choice. so you have \(\dfrac{1}{4}\) equal chance of drawing either of the tiles. For the spinner, you have 5 options and 3 of them are odd. so you have \(\dfrac{3}{5}\) chance of getting a odd number. The combined probability is \(\dfrac{1}{4}*\dfrac{3}{5} =\dfrac{3}{20}=0.15\)